CROATIAN MATHEMATICAL SOCIETY

Bijenička 30, 10 000 Zagreb, Croatia

phone: (385-1)-4555-7201; fax: (385-1)-432-484; e-mail: congress@math.hr

1st CROATIAN MATHEMATICAL CONGRESS

Zagreb, Croatia July 18–20, 1996

FIRST ANNOUNCEMENT

The Croatian Mathematical Society is pleased to announce the First Croatian Mathematical Congress, to be held in Zagreb, Croatia, July 18– 20, 1996. Congress will have a strong international component, and it is open to all areas of mathematics.

Congress activities will start on Thursday morning, July 18, and will end on Saturday afternoon, July 20, 1996. We would like to emphasize that Sunday, July 21, 1996 is also the arrival date for the 2nd European Congress, to be held in nearby Budapest, Hungary.

The chairman of the Scientific Committee is Branko Najman (University of Zagreb). The tentative list of invited speakers includes: J. Azéma (Paris), J. Bertoin (Paris), A. Bourgeat (Saint-Etienne), R.J. Daverman (Knoxville), Z. Janko (Heidelberg), A. Ivić Weiss (York, Canada), A. Mikelić (Lyon), D. Miličić (Salt Lake City), S.J. Patterson (Göttingen), M. Rao (Gainesville, Florida), J.R.M. Sanjurjo (Madrid), K. Veselić (Hagen), M.V. Wickerhauser (St. Louis).

Those interested in receiving additional information should contact us prior to January 1, 1996. All correspondence may be addressed to: Hrvoje Šikić, Department of Mathematics, University of Zagreb, Bijenička 30, 10 000 Zagreb, Croatia.